OpenStudy (anonymous):

what is the solution to the rational equation x/2x+1 + 1/4 = 2/2x+1

OpenStudy (anonymous):

7/2

OpenStudy (anonymous):

sorry -7/2

sam (.sam.):

@khurramshahzad depends, is it \[\frac{x}{2x+1} + \frac{1}{4} = \frac{2}{2x+1}\]

sam (.sam.):

@itsallycatt is your question \[\frac{x}{2x+1} + \frac{1}{4} = \frac{2}{2x+1}\] ?

OpenStudy (anonymous):

yup this is questin

sam (.sam.):

Since you've gone offline, and for other people would like this question solved for reference then

OpenStudy (anonymous):

u r ryt its 7/6 ... i solved it on note book now.. i was stuck in signs .. when doing in my mind withiut notebook thanks for correction